# Quantum Error Correction near the Coding Theoretical Bound

### v2

@@
Abstract
Recent advancements in quantum computing have led to the realization of systems comprising tens of reliable logical qubits, constructed from thousands of noisy physical qubits [1]. However, many of the critical applications that quantum computers aim to solve require quantum computations involving millions or more logical qubits [2, 3]. This necessitates highly efficient quantum error correction capable of handling large numbers of logical qubits. Classical error correction theory is well-developed, with low-density parity-check (LDPC) codes [4] achieving performance limits by encoding large classical bits [kudekar2011threshold]. Despite more than two decades of effort, no efficiently decodable quantum error-correcting code that approaches the hashing bound [6, 7, 8], which is a fundamental lower bound on quantum capacity, had been discovered. Here, we present quantum error-correcting codes constructed from classical LDPC codes that approach the hashing bound while maintaining linear computational complexity in the number of physical qubits. This result establishes a pathway toward realizing large-scale, fault-tolerant quantum computers. By integrating our quantum error correction scheme with devices capable of managing vast numbers of qubits, the prospect of solving critical real-world problems through quantum computation is brought significantly closer.
@@ 1 Introduction
Recent progress in quantum computing has enabled the development of systems with tens of reliable logical qubits, built from thousands of noisy physical qubits [1]. Nonetheless, many essential applications targeted by quantum computers demand quantum computations that involve millions or even more logical qubits [2, 3]. This creates a pressing need for highly efficient quantum error correction methods that can accommodate a vast number of logical qubits.
@@
Protograph codes [9] are LDPC codes whose parity-check matrices are constructed using a protograph matrix composed of permutation matrices as submatrices. A subclass of these codes includes quasi-cyclic (QC) LDPC codes [10], which are constructed using circulant permutation matrices (CPMs), and APM-LDPC (affine permutation matrices) codes [11], which are constructed using APMs. Similar to unstructured LDPC codes, protograph codes are also known for their excellent error correction performance.
Short cycles are particularly known to cause a high error floor. The error floor of LDPC codes refers to a phenomenon where the error rate does not decrease as rapidly as expected with decreasing noise levels, especially in the high-fidelity regime. In computational error correction, extremely low error rates are required, making it critically important to address the error floor phenomenon. It is known that the girth (the length of the shortest cycle) of QC-LDPC matrices is upper-bounded by 12 [10]. However, APM-LDPC matrices can exceed this upper bound on girth [11, 12].
LDPC codes are typically defined over a binary finite field. Certain types of LDPC codes are known to achieve further improvements in error-correcting performance by applying extensions to non-binary finite fields [13], although this increases the computational complexity required for decoding.
@@
CSS codes defined by orthogonal sparse parity-check matrix pairs are referred to as LDPC-CSS codes. Since LDPC-CSS codes are decoded using algorithms similar to those for classical LDPC codes, they are particularly promising for practical application in large-scale quantum computers. Algebraic construction methods allow for the straightforward construction of orthogonal pairs of QC-LDPC matrices [16].
Kasai et al. [17] successfully constructed finite field extended orthogonal non-binary QC-LDPC matrix pairs from the orthogonal QC-LDPC matrix pairs proposed by Hagiwara et al. [16]. To the best of our knowledge, such codes exhibit quantum error-correcting performance that is closest to the hashing bound. However, since the girth of the original QC-LDPC codes is upper-bounded by 12 [10], these codes also exhibited a high error floor.
To date, in previous studies, no quantum error-correcting codes approaching the hashing bound and decodable with computational complexity proportional to the number of physical qubits had been discovered. This study aims to construct LDPC-CSS codes that simultaneously achieve a deep error floor, approach the hashing bound, and are decodable with computational complexity proportional to the number of physical qubits.
@@
Furthermore, we employ a decoding algorithm [18] that decodes X and Z errors simultaneously. In this study, we consider a depolarizing channel where X, Y, and Z errors occur with equal probability, resulting in correlated X/Z errors. In [17], X and Z errors were decoded separately without adopting this decoding method. The proposed algorithm is naturally derived as a sum-product algorithm that marginalizes the posterior probability distributions of X and Z errors under the observed syndrome. Numerical experiments revealed that the error correction performance approaches the hashing bound of a depolarizing channel, exhibiting a deep error floor.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the proposed methods for code construction and decoding. Section 3 presents the error correction performance results of the proposed codes. Further details of these sections are provided in the appendix.
2 Methods
In this section, we briefly describe the construction method and decoding method of the proposed code. For further details, please refer to the appendix.
@@ 2.1 Code construction method
[…]
@@
    Replace the non-zero $\mathbb{F}_{q}$ elements of $H_{\Gamma}$ and $H_{\Delta}$ with companion matrices, which are $\mathbb{F}_{2}$-valued submatrices of size $e\times e$, to construct orthogonal $\mathbb{F}_{2}$-valued matrix pairs $(H_{X},H_{Z})$. Replace 00 with the zero matrix.
Each step is detailed in the appendix. The resulting orthogonal binary matrix pairs $(H_{X},H_{Z})$ both are of size $ePJ\times ePL$. These matrix pairs $(H_{X},H_{Z})$ define the proposed code as orthogonal matrix pairs. Although the proposed code is defined by binary matrix pairs, these matrix pairs can also be interpreted as non-binary $\mathbb{F}_{q}$-valued matrices. Thus, the proposed code aims to achieve high decoding performance by treating it as a non-binary LDPC code during decoding.
@@ 2.2 Decoding method
The sum-product (SP) algorithm is an efficient algorithm for marginalizing multivariate sparsely-factorized functions defined over a large number of variables [20]. Error correction for the proposed codes is performed as follows. Given the observed syndromes $(\underline{s},\underline{t})$, the posterior probability distribution of the errors $(\underline{x},\underline{z})$ are marginalized using the SP algorithm on the sparsely factorized representation. The error estimates $(\hat{\underline{x}},\hat{\underline{z}})$, which maximize the maginal posterior probability distribution, are then determined.
3 Results
First, we randomly constructed $(\underline{f},\underline{g})$ to satisfy the conditions (a), (b), and (c) defined in the appendix. Next, we generated $(\hat{H}_{X},\hat{H}_{Z})$ using the parameters $J=2$, $L\in\{8,10,16\}$, and $P\in\{32,128,1024,8192\}$. These were then extended to $(H_{\Gamma},H_{\Delta})$ through a finite field extension with $e=8$ and subsequently converted to $(H_{X},H_{Z})$ using the companion matrix. We used a $\mathbb{F}_{q=2^{e}}$ with primitive polynomial $1+x^{2}+x^{3}+x^{4}+x^{8}$. You can find these matrices in the file submitted alongside this paper.
The CSS codes defined by $(H_{X},H_{Z})$ are of coding rate $R=1-2J/L$. As a result, the coding rates of the CSS codes constructed above are $R\in\{0.50,0.60,0.75\}$. For the CSS codes defined by $(H_{X},H_{Z})$, the quantum error-correction performance over a depolarizing channel was evaluated. The number of physical qubit is given by $n=ePL$.
@@ 3 Results
In a depolarizing channel, the probabilities of X, Y, and Z errors are each given by $p_{D}/3$. In this case, the marginal probability that either X or Z error occurs is $f_{m}=2p_{D}/3$. The decoding result is defined as successful if the decoding algorithm correctly estimates both error vectors: $\underline{x}=\hat{\underline{x}},\underline{z}=\hat{\underline{z}}$. If the estimation fails for even a single bit of $\underline{x}$ or $\underline{z}$, the decoding is considered failed. The frame error rate (FER) is defined as the ratio of the number of experiments where at least one bit was incorrectly estimated to the total number of experiments. Note that we do not take degenerate errors [21] into consideration. They are treated as decoding failures.
@@
In Figure 1, the decoding performance $(f_{m},\mathrm{FER})$ is plotted for each rate $R$ (from left to right, $R=0.75,0.60,0.50$) by varying the parameter $P$, where the number of physical qubits $n=ePL$ is proportional to $P$. Furthermore, high error-floor was not observed, at least down to an FER of $10^{-4}$. No conventional quantum LDPC codes have been reported where the FER shows sharp threshold and no error floor down to $10^{-4}$ [22].
As $P$ increases, we observe an exponential decrease in the FER. Additionally, a threshold phenomenon is evident, where the FER sharply drops as $P$ grows, particularly when $f_{m}<f_{m}^{*}$ for some critical value $f_{m}^{*}$. We plotted circles at the thresholds $f_{m}^{*}$, which correspond to the points where the curves intersect, for each $R$.
@@
In Figure 2, the pairs of thresholds and coding rates $(f_{m}^{*},R)$ are plotted. There is no conventional quantum error correction that achieves both a deep error floor and a threshold phenomenon [22]. Since conventional quantum error correction does not allow for a well-defined threshold, it is not included in this figure as a comparison. The results show that the proposed codes exhibit decoding performance that approaches the hashing bound.
4 Conclusions
We generalize the quantum error correcting code construction method [17] to general protograph matrices. By performing SP decoding that accounts for the correlation between X and Z errors, we achieved quantum error correction with both deep error floor and sharp threshold. Numerical experiments confirmed that the proposed method achieves error correction approaching the hashing bound, which serves as a benchmark for quantum error correction and represents the lower bound of the quantum capacity.
@@ 4 Conclusions
Figure 1: Decoding performance of the proposed code: $f_{m}$ vs. FER. Decoding performance $(f_{m},\mathrm{FER})$ of the proposed code with $e=8$, $J=2$, $L\in\{8,10,16\}$, $P\in\{32,128,1024,8192\}$, and $R\in\{0.50,0.60,0.75\}$. The black solid line indicates the hashing bound. The code length, that is, the number of physical bits, is given by $n=ePL$. It is evident that the proposed method simultaneously achieves both deep error floors and sharp thresholds.
@@
Figure 2: Decoding performance of the proposed code: $f_{m}^{*}$ vs. $R$. The plot shows $(f_{m}^{*},R)$, where $f_{m}^{*}$ represents the threshold and $R$ denotes the quantum coding rate. It can be observed that all rates approach the hashing bound.
Appendix A Details of code construction
A.1 Construction of $(\hat{H}_{X},\hat{H}_{Z})$
In this section, we describe the construction method for orthogonal parity-check matrix pairs $(\hat{H}_{X},\hat{H}_{Z})$, where each matrix has column weight $J$, row weight $L$, and permutation matrices of size $P$ as submatrices. It is known that non-binary LDPC codes with a column weight $J=2$ exhibit good decoding performance. Therefore, in this paper, the column weight $J$ is consistently set to $J=2$. For a permutation mapping $f:\mathbb{Z}_{P}\to\mathbb{Z}_{P}=\{0,\dots,P-1\}$, we denote by the uppercase $F\in\mathbb{F}_{2}^{P\times P}$, which is the permutation matrix of size $P$ that has a 1 in the $(f(c),c)$ entry.
This construction represents a natural generalization of the method proposed in [16].
  1. (a)
    (Commutativity): The elements of the arrays $\underline{f}$ and $\underline{g}$ are chosen so that they commute with each other. Specifically, they satisfy  [^2] This condition can actually be slightly relaxed as follows: $f_{l-j}$ and $f_{k-l}$ for $20\leq l<L/$ 2 and $0\leq j,k<J$ are commutative.
    $f_{i}\circ g_{j}(x)=g_{j}\circ f_{i}(x)\quad\text{for }0\leq i,j<L/2.$ (7)
    This condition ensures333 For $0\leq j,k<J$, $(\hat{H}_{X}\hat{H}^{\mathsf{T}}_{Z})_{j,k}=\sum_{l=0}^{L/2-1}{\bigl{(}F_{-j+l}G_{k-l}+G_{k-l}F_{-j+l}\bigr{)}}=O$. the orthogonality of $\hat{H}_{X}$ and $\hat{H}_{Z}$. Since a cycle is formed along a specific path444 Two paths consist of a sequence of column sub-matrix indices of length $2L$, denoted as $c_{l}^{1}$ and $c_{l}^{2}$ for $l=0,\cdots,L-1$. For row sub-matrices, the traversal alternates vertically. Specifically, $c_{l}^{1}$ is defined as $l$ when $l$ is even and $L-l$ when $l$ is odd, while $c_{l}^{2}$ is defined as $l$ when $l$ is even and $L-l+1$ when $l$ is odd. traversing all sub-matrices, it can be shown that the girth of corresponding graph is upper-bounded by $2L$. The above conditions serve as sufficient conditions for the girth to achieve this upper bound.
  2. (b)
    (Achieving the girth upper bound $2L$): For $20\leq l,l^{\prime}<L/$ 2, $k\in\{0,\pm 1\}$, and $x\in\{0,\ldots,P-1\}$,
    $f_{l}\circ g_{-l+k}(x)$ $\neq f_{l^{\prime}}\circ g_{-l^{\prime}+k}(x),$ (8)
  3. (c)
By incrementally adding elements of $\underline{f}$ and $\underline{g}$ one by one while satisfying the above three conditions, $\underline{f}$ and $\underline{g}$ can be efficiently constructed. In practice, it is preferable to construct $\underline{f}$ and $\underline{g}$ such that the submatrices are not general PMs but instead CPMs or APMs. Here, we describe the construction using matrices composed of general PMs to clarify the scope of the matrix classes where the above construction method is generally applicable.
For $J=2$, $L=4$, and $P=9$, we provide examples of the above construction method. The arrays $\underline{f}=(x+8,7x+7)$ and $\underline{g}=(x+3,x+6)$, constructed using CPMs, satisfy the conditions described above. Similarly, the sets $\underline{f}=(x+1,x+7)$ and $\underline{g}=(x+1,x+5)$, constructed using CPMs, also satisfy the conditions. In both cases, the girth of the resulting matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ is 8.
A.2 Finite field extension to $(H_{\Gamma},H_{\Delta})$

### v3

@@
[…]
Quantum error correction near the coding theoretical bound
[…]
1 Abstract
Recent progress in quantum computing has enabled systems with tens of reliable logical qubits, built from thousands of noisy physical qubits [1]. However, many impactful applications demand quantum computations with millions of logical qubits [2], necessitating highly scalable quantum error correction. In classical information theory, low-density parity-check (LDPC) codes [3] can approach channel capacity efficiently [4]. Yet, no quantum error-correcting codes with efficient decoding have been shown to approach the hashing bound—a fundamental limit on quantum capacity—despite decades of research [5],[6],[7]. Here, we present quantum LDPC codes that not only approach the hashing bound but also allow decoding with computational cost linear in the number of physical qubits. This breakthrough paves the way for large-scale, fault-tolerant quantum computation. Combined with emerging hardware that manages many qubits, our approach brings quantum solutions to important real-world problems significantly closer to reality.
@@ 2 Introduction
Recent progress in quantum computing has enabled the development of systems with tens of reliable logical qubits, built from thousands of noisy physical qubits [1]. Nonetheless, many essential applications targeted by quantum computers demand quantum computations that involve millions or even more logical qubits [2]. This creates a pressing need for highly efficient quantum error correction methods that can accommodate a vast number of logical qubits.
@@
Protograph codes [8] are LDPC codes whose parity-check matrices are constructed using a protograph matrix composed of permutation matrices (PM) as submatrices. A subclass of these codes includes quasi-cyclic (QC) LDPC codes [9], which are constructed using circulant permutation matrices (CPMs), and APM-LDPC (affine permutation matrices) codes [10], which are constructed using APMs. Similar to unstructublack LDPC codes, protograph codes are also known for their excellent error correction performance.
Short cycles are particularly known to cause a high error floor. The error floor of LDPC codes refers to a phenomenon where the error rate does not decrease as rapidly as expected with decreasing noise levels, especially in the high-fidelity regime. In computational error correction, extremely low error rates are requiblack, making it critically important to address the error floor phenomenon. It is known that the girth (the length of the shortest cycle) of QC-LDPC matrices is upper-bounded by 12 [9]. However, APM-LDPC matrices can exceed this upper bound on girth [10, 11].
LDPC codes are typically defined over a binary field. Certain types of LDPC codes are known to achieve further improvements in error-correcting performance by applying extensions to non-binary finite fields [12], although this increases the computational complexity requiblack for decoding.
@@
CSS codes defined by orthogonal sparse parity-check matrix pairs are referblack to as LDPC-CSS codes. Since LDPC-CSS codes are decoded using algorithms similar to those for classical LDPC codes, they are particularly promising for practical application in large-scale quantum computers. Algebraic construction methods allow for the straightforward construction of orthogonal pairs of QC-LDPC matrices [15].
Kasai et al. [16] successfully constructed finite field extended orthogonal non-binary QC-LDPC matrix pairs from the orthogonal QC-LDPC matrix pairs proposed by Hagiwara et al. [15]. To the best of our knowledge, such codes exhibit quantum error-correcting performance that is closest to the hashing bound. However, since the girth of the original QC-LDPC codes is upper-bounded by 12 [9], these codes also exhibited a high error floor. Furthermore, Hagiwara’s construction method imposes strict constraints on the allowable code lengths; even the subsequent work based on it [16] was limited to code lengths containing large prime factors.
To date, in previous studies, no quantum error-correcting codes approaching the hashing bound and decodable with computational complexity proportional to the number of physical qubits had been discoveblack. This study aims to construct LDPC-CSS codes that simultaneously achieve a error floor down to FER $10^{-4}$, approach the hashing bound, and are decodable with computational complexity proportional to the number of physical qubits.
@@
Furthermore, we employ a efficient decoding algorithm [17] that decodes X and Z errors simultaneously. Although originally designed for binary codes, MacKay’s algorithm is naturally generalized to finite fields in our framework. In this study, we consider a depolarizing channel where X, Y, and Z errors occur with equal probability, resulting in correlated X/Z errors.
@@
In [16], X and Z errors were decoded separately without adopting this decoding method. The proposed algorithm is naturally derived as a sum-product algorithm that marginalizes the posterior probability distributions of X and Z errors under the observed syndrome. Numerical experiments revealed that the error correction performance approaches the hashing bound of a depolarizing channel, with no error floor observed down to FER $10^{-4}$.
@@
The remainder of this paper is organized as follows. In Section 3.1, we describe the proposed methods for code construction and decoding. Section 3.5 presents the error correction performance results of the proposed codes. Further details of these sections are provided in the following section.
3 Results
3.1 Code Construction
As Fossorier [9] pointed out, any $2\times 3$ block of circulant permutation matrices (CPMs) inevitably contains short cycles, limiting the girth of QC-LDPC codes to 12 due to the commutativity of CPMs. To address this, we generalize the construction by [16] to allow arbitrary (not necessarily circulant) permutation matrices (PMs). This increases structural randomness, enhancing decoding performance, and eliminates the need for large prime factors in the code length, a constraint inherent in Hagiwara’s method [15].
3.2 Permutation Matrices
In this section, we describe the construction method for orthogonal parity-check matrix pairs $(\hat{H}_{X},\hat{H}_{Z})$, where each matrix has column weight $J$, row weight $L$, and permutation matrices of size $P$ as submatrices. It is known that non-binary LDPC codes with a column weight $J=2$ exhibit good decoding performance. Therefore, in this paper, the column weight $J$ is consistently set to $J=2$. For a permutation mapping $f:\mathbb{Z}_{P}\to\mathbb{Z}_{P}=\{0,\dots,P-1\}$, we denote by the uppercase $F\in\mathbb{F}_{2}^{P\times P}$, which is the permutation matrix of size $P$ that has a 1 in the $(f(c),c)$ entry. The set of such functions is denoted by $\mathcal{F}_{P}$.
Let $P=4$ and consider the permutation $f(x)=3x+2\mod 4$ over $\mathbb{Z}_{4}$. Then, $f(0)=2,\quad f(1)=1,\quad f(2)=0,\quad f(3)=3.$ The corresponding permutation matrix $F\in\mathbb{F}_{2}^{4\times 4}$ has a $1$ in position $(f(c),c)$ and is given by
$F=\begin{bmatrix}0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\end{bmatrix}$.
Here, $F_{i}$ and $G_{i}$ for $i=0,1,\ldots,L/2-1$ are the permutation matrices corresponding to the permutations $f_{i}$ and $g_{i}$, respectively. This construction represents a natural generalization of the method proposed in [15].
  1. (a)
    (Commutativity): The elements of the arrays $\underline{f}$ and $\underline{g}$ are chosen so that they commute with each other. Specifically, they satisfy
    $f_{i}\circ g_{j}(x)=g_{j}\circ f_{i}(x)\quad\text{for }0\leq i,j<L/2.$ (7)
    In fact, this condition can be slightly relaxed. It suffices that $f_{l-j}$ and $g_{k-l}$ commute for $0\leq l<L/2$ and $0\leq j,k<J$. This commutativity condition guarantees the orthogonality of $\hat{H}_{X}$ and $\hat{H}_{Z}$, as it ensures that for $0\leq j,k<J$,
    $(\hat{H}_{X}\hat{H}^{\mathsf{T}}_{Z})_{j,k}=\sum_{l=0}^{L/2-1}{\bigl(F_{-j+l}G_{k-l}+G_{k-l}F_{-j+l}\bigr)}=O.$ (8)
    Since a cycle is formed along a specific path traversing all sub-matrices, it can be shown that the girth of corresponding graph is upper-bounded by $2L$. Two paths consist of a sequence of column sub-matrix indices of length $2L$, denoted as $c_{l}^{1}$ and $c_{l}^{2}$ for $l=0,\cdots,L-1$. For row sub-matrices, the traversal alternates vertically. Specifically, $c_{l}^{1}$ is defined as $l$ when $l$ is even and LlL-l when $l$ is odd, while $c_{l}^{2}$ is defined as $l$ when $l$ is even and Ll+1L-l+1 when $l$ is odd.
  2. (b)
    (Achieving the girth upper bound $2L$): For $0\leq l,l^{\prime}<L/2$, $k\in\{0,\pm 1\}$, and $x\in\{0,\ldots,P-1\}$,
    $f_{l}\circ g_{-l+k}(x)$ $\neq f_{l^{\prime}}\circ g_{-l^{\prime}+k}(x),$ (9)
  3. (c)
The sets $\underline{f}$ and $\underline{g}$ can be efficiently constructed by sequentially adding their elements while ensuring that the three conditions described above are satisfied at each step. The concrete procedure is provided in Algorithm 1. While general PMs are used in this description to clarify the general applicability of the construction method, in practice it is often preferable to employ CPMs or APMs.
Algorithm 1 Sequential Construction of $\underline{f}=(f_{0},\ldots,f_{L/2-1}),\underline{g}=(g_{0},\ldots,g_{L/2-1})$
1:Initialize empty list $\mathcal{S}\leftarrow[\ ]$
2:for $i=0$ to L/21L/2-1 do
3:  repeat
4:   Randomly generate a candidate $f_{i}$ from $\mathcal{F}_{P}$
5:   Temporarily set $\mathcal{S}^{\prime}\leftarrow\mathcal{S}\cup\{f_{i}\}$
6:   if conditions (a), (b), and (c) are satisfied for the current subset $\mathcal{S}^{\prime}$ then
7:     Accept $f_{i}$: $\mathcal{S}\leftarrow\mathcal{S}^{\prime}$
8:     break
9:   end if
10:  until a valid $f_{i}$ is found
11:  repeat
12:   Randomly generate a candidate $g_{i}$ from $\mathcal{F}_{P}$
13:   Temporarily set $\mathcal{S}^{\prime}\leftarrow\mathcal{S}\cup\{g_{i}\}$
14:   if conditions (a), (b), and (c) are satisfied for the current subset $\mathcal{S}^{\prime}$ then
15:     Accept $g_{i}$: $\mathcal{S}\leftarrow\mathcal{S}^{\prime}$
16:     break
17:   end if
18:  until a valid $g_{i}$ is found
19:end for
20:return $\mathcal{S}$
For $J=2$, $L=4$, and $P=12$, we present an example of the above construction method. The function arrays $\underline{f}=(5x+4,\ 5x+8)$ and $\underline{g}=(7x+6,\ 7x+9)$, constructed using APMs, satisfy the requiblack conditions. In this case, the girth of the resulting parity-check matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ is 8. Figure 1 shows the matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ derived from $\underline{f}$ and $\underline{g}$, where black cells indicate ones and white cells indicate zeros.
Figure 1: An example of $\hat{H}_{X}$ and $\hat{H}_{Z}$
@@ 3.3 Code Construction Method
In this section, we describe how to construct the matrices $H_{X}$ and $H_{Z}$ from the previously constructed $\hat{H}_{X}$ and $\hat{H}_{Z}$. As the procedure largely overlaps with the method presented in [16], we defer the detailed explanation to the following section.
The proposed quantum error correction employs CSS codes defined by orthogonal $\mathbb{F}_{2}$-valued matrix pairs $(H_{X},H_{Z})$, where $\mathbb{F}_{2}=\{0,1\}$. These matrix pairs are obtained through the following three steps:
@@
    Replace the non-zero $\mathbb{F}_{q}$ elements of $H_{\Gamma}$ and $H_{\Delta}$ with companion matrices, which are $\mathbb{F}_{2}$-valued submatrices of size $e\times e$, to construct orthogonal $\mathbb{F}_{2}$-valued matrix pairs $(H_{X},H_{Z})$. Replace 0 with the zero matrix.
Each step is detailed in the following section. The resulting orthogonal binary matrix pairs $(H_{X},H_{Z})$ both are of size $ePJ\times ePL$. These matrix pairs $(H_{X},H_{Z})$ define the proposed code as orthogonal matrix pairs. Although the proposed code is defined by binary matrix pairs, these matrix pairs can also be interpreted as non-binary $\mathbb{F}_{q}$-valued matrices. Thus, the proposed code aims to achieve high decoding performance by treating it as a non-binary LDPC code during decoding.
@@ 3.4 Decoding method
[…]
@@
In a depolarizing channel, the probabilities of X, Y, and Z errors are each given by $p_{D}/3$. In this case, the marginal probability that either X or Z error occurs is $f_{m}=2p_{D}/3$. The decoding result is defined as successful if the decoding algorithm correctly estimates both error vectors: $\underline{x}=\hat{\underline{x}},\underline{z}=\hat{\underline{z}}$. If the estimation fails for even a single bit of $\underline{x}$ or $\underline{z}$, the decoding is consideblack failed. The frame error rate (FER) is defined as the ratio of the number of experiments where at least one bit was incorrectly estimated to the total number of experiments.
Note that we do not take degenerate errors [21] into consideration; they are treated as decoding failures. Properly addressing degeneracy is generally consideblack essential for approaching the Hashing bound. To close the remaining gap to the Hashing limit, it will be necessary to explicitly handle degeneracy. However, a rigorous treatment of degeneracy is beyond the scope of this work and is left for future study.
While our decoder does not explicitly account for degeneracy, we adopt a strict criterion that treats any mismatch from the true noise—regardless of logical equivalence—as a failure. As such, the reported FER should be interpreted as conservative estimates. For the purposes of this study, the omission of degeneracy is not expected to affect asymptotic behaviors such as the waterfall region, although it may impact the error floor.
3.5 Numerical Results
First, we randomly constructed $(\underline{f},\underline{g})$ to satisfy the conditions (a), (b), and (c) defined in the following section. Next, we generated $(\hat{H}_{X},\hat{H}_{Z})$ using the parameters $J=2$, $L\in\{8,10,16\}$, and $P\in\{32,128,1024,8192\}$. These were then extended to $(H_{\Gamma},H_{\Delta})$ through a finite field extension with $e=8$ and subsequently converted to $(H_{X},H_{Z})$ using the companion matrix. We used a $\mathbb{F}_{q=2^{e}}$ with primitive polynomial $1+x^{2}+x^{3}+x^{4}+x^{8}$.
The CSS codes defined by $(H_{X},H_{Z})$ are of coding rate R=12J/LR=1-2J/L. As a result, the coding rates of the CSS codes constructed above are $R\in\{0.50,0.60,0.75\}$. For the CSS codes defined by $(H_{X},H_{Z})$, the quantum error-correction performance over a depolarizing channel was evaluated. The number of physical qubit is given by $n=ePL$. The parameters of the codes used in the experiments are listed in Table 1. Here, $n$ denotes the number of physical qubits, $k$ is the number of logical bits, and $\overline{d}$ represents an upper bound on the minimum distance. Since determining the exact minimum distance is computationally difficult, we estimated an upper bound using the following approach.
Table 1: Parameters of the proposed codes • and the code $\circ$ [16] that was previously consideblack to be the closest to the hashing bound in [16].
$n$ $k$ $\overline{d}$ $P$ $J$ $L$ $R$ $e$
8192 4096 10 128 2 8 0.50 8
65536 32768 11 1024 2 8 0.50 8
524288 262144 10 8192 2 8 0.50 8
$\circ$ 8768 4384 - 137 2 8 0.50 8
2560 1536 9 32 2 10 0.60 8
10240 6144 10 128 2 10 0.60 8
81920 49152 9 1024 2 10 0.60 8
[…]
4096 3072 9 32 2 16 0.75 8
16384 12288 8 128 2 16 0.75 8
131072 98304 9 1024 2 16 0.75 8
@@ 3.5 Numerical Results
In Figure 2, the decoding performance $(f_{m},\mathrm{FER})$ is plotted for each rate $R$ (from left to right, $R=0.75,0.60,0.50$) by varying the parameter $P$, where the number of physical qubits $n=ePL$ is proportional to $P$. Furthermore, high error-floor was not observed, at least down to an FER of $10^{-4}$. No conventional quantum LDPC codes have been reported where the FER shows error floors down to FER $10^{-4}$ and waterfall near the hashing bound [22, 23]. We also compare our results with those of Kasai et al. (2011), which were obtained under similar parameters. As indicated by the gray arrow, a significant performance improvement is observed.
Although this paper focuses on the case $L\geq 8$, recent results [24] have shown that relatively high error floors can emerge at large block lengths when $L=6$. To address this issue, our recent work [24] demonstrates that degeneracy-aware post-processing, which specifically targets stopping sets, can significantly blackuce the FER in the low-noise regime.
For codes with row weight $L\geq 8$, no error floor was observed down to a FER of $10^{-4}$. Nevertheless, the possibility of an error floor emerging below this level cannot be ruled out. Further investigation involving deeper simulations will be necessary to explore the low-FER regime and to gain a more complete understanding of the limitations of the proposed code constructions and the decoding strategy.
@@
Figure 3 plots $(f_{m}^{*},R)$ where $f_{m}^{*}$ denotes the value of $f_{m}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. There is no conventional quantum error correction that achieves both an error floor down to FER $10^{-4}$ and steep waterfall near the hashing bound [23]. The results show that the proposed codes exhibit decoding performance that approaches the hashing bound.
@@
Figure 2: Decoding performance of the proposed code: $f_{m}$ vs. FER. Decoding performance $(f_{m},\mathrm{FER})$ of the proposed code with $e=8$, $J=2$, $L\in\{8,10,16\}$, $P\in\{32,128,1024,8192\}$, and $R\in\{0.50,0.60,0.75\}$. The black solid line indicates the hashing bound. The code length, that is, the number of physical bits, is given by $n=ePL$. It is evident that the proposed method simultaneously achieves both error floors down to FER $10^{-4}$ and waterfall near the hashing bound. The gray dashed curve represents a rate-1/2 code [16] that was previously consideblack to be the closest to the hashing bound [22].
@@
Figure 3: Decoding performance of the proposed code: $f_{m}$ vs. $R$. The plot shows the pairs $(f_{m}^{*},R)$, where $f_{m}^{*}$ denotes the value of $f_{m}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. The filled circles represent the results of the proposed method, while the empty circles correspond to the method defined over the same finite field as proposed in [16].
4 Discussion
We generalize the quantum error correcting code construction method [16] to general protograph matrices. By performing SP decoding that accounts for the correlation between X and Z errors, we achieved quantum error correction with both error floors down to FER $10^{-4}$ and waterfall near the hashing bound. Numerical experiments confirmed that the proposed method achieves error correction approaching the hashing bound, which serves as a benchmark for quantum error correction and represents the lower bound of the quantum capacity.
5 Methods
5.1 Finite field extension to $(H_{\Gamma},H_{\Delta})$